OpenStudy (anonymous):

System of equations an example on how to do it. (Since many people here are asking similar questions, I will show you how to do at least one of these problems.) [Elimination method]

OpenStudy (anonymous):

Oh, I was making an example. lol

OpenStudy (anonymous):

Ex. 2x-y+4z = 17 3x+4y-z = -1 4x+3y+2z = 11

OpenStudy (yanasidlinskiy):

Finally at least someone explained!:D Thank you!!!!

OpenStudy (anonymous):

@geekfromthefutur this noble guy is helping people learn it here

OpenStudy (anonymous):

First thing, we'll do is label the equations to make it easier for us to know where what is. 1. 2x-y+4z=17 2. 3x+4y-z=-1 3. 4x+3y+2z=11 Now we'll pick a variable to eliminate, it's arbitrary, so with that said, lets pick z. 1. 2x-y+4z = 17 2. 3x+4y-z = -1 we want to make the variable have the same number but opposite sign, so we can eliminate it. So we can take 2 and multiply the whole equation by 4 --> 4(3x+4y-z=-1) which would make our new equation 12x+16y-4z = -4. We will call this 2a. Now back to 1 and 2, we can now add these equations together. 1. 2x-y+4z = 17 2a. 12x+16y-4z = -4 Adding these we get 14x+15y=13, we will call this equation 4. 4. 14x+15y=13 Now we have made an easier equation with only two variables, but to continue to eliminate one of these variables we will have to make another. So this time we'll take equations 2 and 3 and repeat the process again. 2. 3x+4y-z=-1 3. 4x+3y+2z = 11 Multiply equation 1 by 2, so 2(3x+4y-z = -1), our new equation will be 6x+8y-2z = -2 we will call this 2b. Now we can add these two equations as before, 2b. 6x+8y-2z=-2 3. 4x+3y+2z=11 Adding these equations we get 10x+11y=9, we call this 5. 5. 10x+11y= 9 Using equations 4 and 5, we can now find the variables x and y. 4. 14x+15y=13 5. 10x+11y = 9 We'll find y first, to do this, we want to eliminate x. Multiply equation 4 by -5 --> -5(14x+15y=13) --> -70x-75y=-65 we'll call this 4a. Multiply equation 5 by 7 --> 7(10x+11y=9) --> 70x+77y = 63 we'll call this 5a. Adding these two equations together 4a. 70x-75y=-65 5a. 70x +77y = 63 We get 2y = -2, dividing both sides by 2, y = -1. Now we can pick equations 4 or 5 to substitute y in, we'll pick 5. 5. 10x+11y = 9 Substituting y = -1 we get 10x+11(-1) = 9 => x = 2. Now we have x = 2, and y = -1, now the last part is just to find z. To do this we pick one of the original functions and substitute x = 2, and y = -1 for it. Here we'll pick equation 2 (it doesn't matter which one you pick). 2. 3x+4y-z=-1 Now we substitute the variables x and y. 3(2)+4(-1)-z = -1 6-4-z = -1 -z = -3 divide both sides by -1. We get z = 3. Final answer = (2,-1,3). It looks a bit complicated but when you start doing it, you'll see it's actually very easy.

OpenStudy (anonymous):

Just ask any questions if you don't get a specific part.

OpenStudy (anonymous):

@CountryDoe13 @ilovehim121511 This might help you.

OpenStudy (somy):

lol thank god when i used to do system of equations or whatever u call it it wasn't this big lol

OpenStudy (somy):

uh well i just remembered that i still do it in physics <.<

OpenStudy (anonymous):

It's actually not that big haha, I was just trying to make the steps easier, by symbolizing each part, so it would be easier for the people to follow it.

OpenStudy (anonymous):

I feel as if I need to make legit tutorials for all of high school math lol.

OpenStudy (somy):

3 equations seems pretty big tho thank god i know this stuff >.> i'd die in physics if i didn't

OpenStudy (somy):

lol go for it u r a Batmaaan u can do it lmao Good job tho!!!!

OpenStudy (anonymous):

Lol, maybe later.

OpenStudy (somy):

@Jaleena take a look at this @iambatman explains it pretty good

sammixboo (sammixboo):

Explanations *v(=u=*v) Yay

OpenStudy (anonymous):

Who here is taking edgenuity

OpenStudy (anonymous):

'Cause he's batman

OpenStudy (amistre64):

elimination is similar to the matrix solution

OpenStudy (anonymous):

Elimination is similar to shooting a deer.

OpenStudy (aum):

People should first master solving two equations with two unknowns. Then three equations with three unknowns can be reduced to two equations with two unknowns by picking one variable to eliminate between equations (1) and (2) and between (2) and (3) or (1) and (3). Then all three unknowns can be solved.